OpenStudy (anonymous):

solve this___ Simplify i31

OpenStudy (mertsj):

It is a prime number

OpenStudy (anonymous):

Simplify square root of negative 72

OpenStudy (mertsj):

\[\sqrt{-72}=\sqrt{36(2)(-1)}=\sqrt{36}\sqrt{2}\sqrt{-1}=6\sqrt{2}i\]